# CMQCIC-Bench: A Chinese Benchmark for Evaluating Large Language Models in Medical Quality Control Indicator Calculation

### v1

@@ Abstract
Medical quality control indicators are essential to assess the qualifications of healthcare institutions for medical services. With the impressive performance of large language models (LLMs) like GPT-4 in the medical field, leveraging these technologies for the Medical Quality Control Indicator Calculation (MQCIC) presents a promising approach. In this work, (1) we introduce a real-world task MQCIC and propose an open-source Chinese electronic medical records (EMRs)-based dataset (CMQCIC-Bench) comprising 785 instances and 76 indicators. (2) We propose a semi-automatic method to enhance the rule representation. Then we propose the Clinical Facts-based Inferential Rule (CF-IR) method that disentangles the clinical fact verification and inferential rule reasoning actions. (3) We conduct comprehensive experiments on 20 representative LLMs, covering general and medical models. Our findings reveal that CF-IR outperforms Chain-of-Thought methods in MQCIC tasks. (4) We conduct an error analysis and investigate the capabilities of clinical fact verification and inferential rule reasoning, providing insights to improve performance in the MQCIC further. The dataset and code is available in this repo 111https://anonymous.4open.science/r/C-MQCIC-1151.
@@
Guangya Yu1, Yanhao Li1, Zongying Jiang1, Yuxiong Jin1, Li Dai1, Yupian Lin1, Ruihui Hou1, Weiyan Zhang1, Yongqi Fan1, Qi Ye1, Jingping Liu1, Tong Ruan1 1School of Information Science and Engineering, East China University of Science and Technology, Shanghai, China guangyayu@mail.ecust.edu.cn
@@ 1 Introduction
Medical quality control indicators play an essential role in assessing the performance of healthcare institutions Øvretveit (2001); Wang et al. (2018); Anderson et al. (2017). Recently, Large Language Models (LLMs) like GPT-4 Achiam et al. (2023) have shown promising capabilities in the medical domain. These include applications such as diagnostic reasoning Dou et al. (2024), clinical note generation Yang et al. (2023a), and automated clinical assessment GU et al. (2024). Such capabilities may also prove effective in Medical Quality Control Indicator Calculation (MQCIC).
@@ 3 Dataset Construction
In this section, we construct a dataset CMQCIC-Bench for the MQCIC task. The main content includes the data collection of indicators and patient notes, data annotation, and data characteristics.
@@ 3.2 Data Annotation
Specifically, the annotation process uses the following three-step pipelines. (1) Clinical fact extraction. We leverage GPT-4o to extract the original information from EMRs without any modification, then reason based on the context to verify the clinical fact. The clinical facts contain GCS scores, lab exams, medications, etc. (2) Answer and explanation generation. Given the obtained facts, for each instance, we leverage GPT-4o to generate the step-by-step explanation through logical reasoning and a final answer . (3) Data quality control. Finally, with the guidance of medical experts, annotators are required to check the answer in three facets: fact extraction, logical reasoning, and consistency. Fact extraction and logical reasoning verify the accuracy of the first two stages, while consistency ensures alignment within the patient notes to exclude low-quality cases. In the end, we curated 785 instances for CMQCIC-Bench, as shown in Figure 2, each instance consists of a Patient Note , a Question , a step-by-step , and the final answer . With the same process, we constructed a CMQCIC-Private dataset derived from patient notes of top-tier tertiary hospitals in China. Ethics committees and experts have rigorously de-identified these data to ensure no privacy leakage risk.
@@ 4 Method
To address ambiguity in indicator rules, we propose a semi-automatic method that decomposes them into transparent templated clinical facts and logical rules. We then introduce the Clinical Fact-based Inferential Rule (CF-IR) reasoning pattern for inference. As illustrated in Figure 4, the method comprises two key components: Rule Representation Enhancement and CF-IR.
@@ 4.1 Rule Representation Enhancement
Step 1: Knowledge Enhancement. We leverage GPT-4o to recall relevant information instead of collecting additional medical books and guidelines.
@@
Step 3: Clinical Fact Templatization. GPT-4o further extracts the clinical facts involved in the logical rules, which are independent. Each clinical fact with a supposed answer set such as True/False, numerical value, etc.
In the end, we enlisted human experts to review the enhanced rules for the 76 indicators. We also explore the automatic performance in Section 5.2.
@@ 4.2 Clinical Fact-based Inferential Rule Reasoning
(1) Clinical Fact Verification. Before engaging in clinical reasoning, it is crucial to obtain accurate clinical informationWang et al. . However, extracting clinical facts and verifying them through reasoning in a long context with noises of over 380.41 tokens is quite challenging. This includes identifying synonyms, linking symptoms to facts (e.g., high GCS indicates better consciousness), understanding medications (e.g., dual antiplatelet therapy), and analyzing surgical indications, all requiring medical knowledge and clinical reasoning.
@@
Human | 95.00 |
(2) Inferential Rule Reasoning. Reasoning can be categorized in multiple ways Yu et al. (2024). In this study, we specifically define ’Inferential Rule Reasoning’ as the capability to deduce a final conclusion by applying logical rules to multiple clinical facts. RuleBench evaluates inferential rule-following abilities Sun et al. (2024) in scenarios such as math QA, and relation extraction, while LegalBench Guha et al. (2024) focuses on the legal domain. In contrast, our dataset evaluates the capability within clinical scenarios, requiring LLMs to have relevant medical knowledge and a deep understanding of complex rules.
@@ 5.1 Settings
Models. We select a representative set of 14 general large language models including GPT-4555gpt-4o-2024-08-06 Achiam et al. (2023), Qwen2.5 series Yang et al. (2024a), Internlm2.5 series Cai et al. (2024), Llama3.1 series Dubey et al. (2024), MiniCPM3-4B Hu et al. (2024),as well as 6 medical large language models, comprising the Apollo series Zheng et al. (2025); Wang et al. (2024b), and HuatuoGPT2 series Chen et al. (2023). We omit the "chat" and "instruct" in Table 2.
@@
Methods. We assess six prompt-based methodologies: (1) Standard Prompt, utilizing solely the original rules and patient notes; (2) Zero-Shot CoT Kojima et al. (2022), enhanced with the directive "Let’s think step by step"; (3) We leverage the CF-IR method to derive the answer; (4) Specifically, we examine the One-shot CoT Wei et al. (2022); (5) Similar to Chain-of-Logic Servantez et al. (2024b), we introduce a setting, (ACF-IR), designed to enable LLMs to decompose rules and perform reasoning based on them independently; (6) We set a One-Shot CF-IR. For each indicator, we selected an example outside of CMQCIC-Bench. The outputs for these examples were generated by GPT-4o in a zero-shot setting and were carefully reviewed and annotated by human evaluators.
@@ 5.2 Main Results
6 Emperical Analysis and Discussion
@@ 6.1 Error Analysis
Building on prior work Khandekar et al. (2025), we employ DeepSeek to classify error types by comparing LLM outputs with ground truth in CMQCIC-Bench, facilitating a granular error analysis across LLMs. Since incorrect clinical facts can propagate and affect inferential rule reasoning, we focus on identifying the earliest error type. A manual review of 200 randomly sampled DeepSeek-annotated errors confirmed an 87% accuracy, validating our approach for analyzing error types in all CF-IR responses. As shown in Table 4, providing demonstrations reduces Type A and B errors, highlighting the value of exemplars. While CF-IR does not mitigate clinical fact verification errors, it significantly improves reasoning accuracy due to its structured logical framework. Further details are available in Appendix F.1 (Tables 9 and 10).
@@ 6.3 Analysis on Clinical Fact Verification and Inferential Rule Abilities
While the CF-IR method enhances inference performance, we further investigate its two core capabilities: Clinical Fact Verification and Inferential Rule Reasoning. For Clinical Fact Verification, we define the input as <Patient Note, Templated Clinical Fact, Question> and the output as <Reasoning, Final Answer>, evaluated using Fact Faithfulness and Fact Correctness. Unlike step-wise reasoning, we test each fact independently to avoid contextual interference. For Inferential Rule Reasoning, providing verified facts as input to minimize errors, the input is <Verified Clinical Facts, Logical Rules>, and the output is <Explanation, Final Answer>, evaluated using labels like ’True/Yes’ or ’False/No’ for both natural (NL) and symbolic (SY) languages. All experiments are conducted in a zero-shot setting. Additional results are available in Table 11 in Appendix F.2.
As shown in Figures 5 we find that: (1) Both Qwen and Apollo exhibit performance-scale correlations across capabilities. (2) Fact verification performance significantly declines, consistent with Table 5. (3) For inferential reasoning, Qwen performs comparably in natural and symbolic settings, while Apollo shows stronger natural language robustness. (4) With correct facts, Qwen surpasses previous best results (standard, CoT, CF-IR) in zero-shot settings, whereas Apollo underperforms, likely due to Qwen’s extensive logical reasoning training. See Appendix F.2 for additional results.
@@ 6.4 Benefit of fine-tuning
We fine-tuned the Qwen2.5-3B-Instruct model using LoRA Hu et al. (2021) for 3 epochs on the CMQCIC-Bench, with evaluation on CMQCIC-Private. The data format follows ACF-IR: Input: <Instruction, Patient Note, Rule>; Output: <Knowledge, Templated Clinical Facts, Logical Rules, Clinical Fact Verification, Inferential Rule Reasoning, Final Answer>. As depicted in Figure 6, the fine-tuned 3B model achieves comparable or superior performance to the 7B model in real-world scenarios. In zero-shot settings, it demonstrates significant improvements across all methods, with gains of 6.7, 8.3, 14.3, and 13.7, confirming the feasibility of distillation rule enhancement for smaller models. More details are provided in Appendix D.
@@ 7.1 Rule-based LLM reasoning
While reasoning demonstrated a fundamental capability of LLM on applications Li et al. (2024), there are many research such as CoT Wei et al. (2022), CoT-sc Wang et al. (2023b), ToT Yao et al. (2024), etc. However, there are more attention on rule-enhanced methods Yang et al. (2023b); Sun et al. (2023); Mu et al. (2023). Reasoning based on facts and deriving answers from logical rules is referred to as inferential rule following ability Sun et al. (2024). Leveraging such ability that integrating explicit rules with LLMs has gained significant attention. For instance, Servantez et al. (2024b) utilized the IRAC framework to tackle legal tasks with LLMs, emphasizing the application of legal rules. Additionally, Wang et al. (2024a) proposed a neurosymbolic framework for multi-step rule application. Despite the current limitations of LLMs in rule-based reasoning Yang et al. (2024c), our work demonstrates that such rule-based reasoning outperforms CoT reasoning in the MQCIC task.
@@ 8 Conclusion
Appendix A CMQCIC-Bench Additional Detail
For each indicator rule question, we provide a list of facts that should be extracted. The prompt for clinical fact extraction is shown in Figure 7. The main fields are in the data.json 8 and indicator.jsonl files 7.
Appendix B Models
@@
Appendix C Inference Cost
As shown in Table 6, we report the inference time of Qwen, which indicates that the CF-IR method can reduce the inference time in a zero-shot setting, while providing a demonstration that may increase the cost in a one-shot setting. Thus, a possible approach for practical application is leveraging strong LLM like GPT-4o to enhance the rule and utilizing the lightweight LLM during inference.
@@ Appendix C Inference Cost
Appendix D Training Details
@@
Appendix E Prompt
Here are the zero-shot prompt templates for the rule representation enhancement, clinical fact-based inferential rule reasoning method, and other prompt-based methods we used in this paper.
Prompt for Rule Enhancement. As mentioned in Section 4, we leverage LLMs to transform rules through three steps: Knowledge Enhancement, Rule Decomposition, and Clinical Fact Templazation. As shown in Figure 8, 9 and 10, we leverage GPT-4o first to generate the relative knowledge, logical rules, and templated clinical facts separately.
Prompt for Different Methods. The detail prompt of different method as shown in Figure 11, Figure 12 and Figure 13.
Appendix F Additional Analysis Details
F.1 Error Analysis Details
As shown in Table 9 and Table 10 we observed: (1) Although the one-shot method reduces errors across both types in both approaches, the CoT method still results in more Type B errors, which may be due to the differing reasoning paths in the examples. (2) The CF-IR method effectively reduces Type B errors, but when it comes to Type A errors, the issue seems to be more related to the model’s intrinsic capabilities, which our method has not been able to effectively enhance or activate.
F.2 Clinical Fact Verification and Inferential Rule

### v2

@@ Abstract
Medical quality control indicators are essential to assess the qualifications of healthcare institutions for medical services. With the impressive performance of large language models (LLMs) like GPT-4 in the medical field, leveraging these technologies for the Medical Quality Control Indicator Calculation (MQCIC) presents a promising approach. In this work, (1) we introduce a real-world task MQCIC and propose an open-source Chinese electronic medical records (EMRs)-based dataset (CMQCIC-Bench) comprising 785 instances and 76 indicators. (2) We propose a semi-automatic method to enhance the rule representation. Then we propose the Clinical Facts-based Inferential Rule (CF-IR) method that disentangles the clinical fact verification and inferential rule reasoning actions. (3) We conduct comprehensive experiments on 20 representative LLMs, covering general and medical models. Our findings reveal that CF-IR outperforms Chain-of-Thought methods in MQCIC tasks. (4) We conduct an error analysis and investigate the capabilities of clinical fact verification and inferential rule reasoning, providing insights to improve performance in the MQCIC further. The dataset and code is available in this repository 111https://github.com/YuY-2001/C-MQCIC.
@@
Guangya Yu, Yanhao Li, Zongying Jiang, Yuxiong Jin, Li Dai, Yupian Lin, Ruihui Hou, Weiyan Zhang, Yongqi Fan, Qi Ye††thanks: Corresponding authors., Jingping Liu, Tong Ruan11footnotemark: 1 School of Information Science and Engineering, East China University of Science and Technology, Shanghai, China guangyayu@mail.ecust.edu.cn, {yeh_qi1125,ruantong}@ecust.edu.cn
@@ 1 Introduction
Medical quality control indicators play an essential role in assessing the performance of healthcare institutions Øvretveit (2001); Wang et al. (2018); Anderson et al. (2017). Recently, Large Language Models (LLMs) like GPT-4 Achiam et al. (2023) have shown promising capabilities in the medical domain. These include applications such as diagnostic reasoning Dou et al. (2024), clinical note generation Yang et al. (2023a), and automated clinical assessment GU et al. (2024). Such capabilities also prove effective in Medical Quality Control Indicator Calculation (MQCIC) Ye et al. (2025).
@@ 3 Dataset Construction
In this section, we construct a dataset, CMQCIC-Bench, for the MQCIC task. The main content includes the data collection of indicators and patient notes, data annotation, and data characteristics.
@@ 3.2 Data Annotation
Specifically, the annotation process uses the following three-step pipeline. (1) Clinical fact extraction. We leverage GPT-4o to extract the original information from EMRs without any modification, then reason based on the context to verify the clinical fact. The clinical facts contain GCS scores, lab exams, medications, etc. (2) Answer and explanation generation. Given the obtained facts, for each instance, we leverage GPT-4o to generate the step-by-step explanation through logical reasoning and a final answer . (3) Data quality control. Finally, with the guidance of medical experts, annotators are required to check the answer in three facets: fact extraction, logical reasoning, and consistency. Fact extraction and logical reasoning verify the accuracy of the first two stages, while consistency ensures alignment within the patient notes to exclude low-quality cases. In the end, we curated 785 instances for CMQCIC-Bench, as shown in Figure 2, each instance consists of a Patient Note , a Question , a step-by-step , and the final answer . With the same process, we constructed a CMQCIC-Private dataset derived from patient notes of top-tier tertiary hospitals in China. Ethics committees and experts have rigorously de-identified these data to ensure no privacy leakage risk.
@@ 4 Method
As shown in Figure 1, the indicators constructed by experts are quite vague, and the underlying medical knowledge can directly affect the implementation of the rules. Therefore, we propose a semi-automatic method that decomposes them into transparent, templated clinical facts and logical rules. We then introduce the Clinical Fact-based Inferential Rule (CF-IR) reasoning pattern for inference. As illustrated in Figure 4, the method comprises two key components: Rule Representation Enhancement and CF-IR. Additionally, we introduce Automatic CF-IR (ACF-IR) to explore the automatic performance of rule representation enhancement.
@@ 4.1 Rule Representation Enhancement
Step 1: Knowledge Enhancement. We leverage GPT-4o to recall relevant information instead of collecting additional medical books and guidelines. It aims to resolve rule ambiguities arising from the lack of knowledge. For example, the high and low scores of GCS and NIHSS have different meanings, and splitting the rules based solely on semantic information could result in incorrect logical rules.
@@
Step 3: Clinical Fact Templatization. GPT-4o further extracts the clinical facts involved in the logical rules, which are independent. Each clinical fact with a supposed answer set, such as True/False, numerical value unit, etc.
In the end, we enlisted human experts to review the enhanced rules for the 76 indicators.
@@ 4.2 Clinical Fact-based Inferential Rule Reasoning
(1) Clinical Fact Verification. Before engaging in clinical reasoning, it is crucial to obtain accurate clinical informationWang et al. (2024a). However, extracting clinical facts and verifying them through reasoning in a long context with noises of over 380.41 tokens is quite challenging. This includes identifying synonyms, linking symptoms to facts (e.g., high GCS indicates better consciousness), understanding medications (e.g., dual antiplatelet therapy), and analyzing surgical indications, all requiring medical knowledge and clinical reasoning.
@@
| Human | 95.00 | ||||||
(2) Inferential Rule Reasoning. Reasoning can be categorized in multiple ways Yu et al. (2024). In this study, we specifically define ’Inferential Rule Reasoning’ as the capability to deduce a final conclusion by applying logical rules to multiple clinical facts. Specifically, for each templated clinical fact, the LLM first performs targeted information extraction and verification. Subsequently, it applies explicit logical rules to conduct structured reasoning based on the verified clinical facts.
@@ 5.1 Settings
Models. We select a representative set of 14 general large language models including GPT-4555gpt-4o-2024-08-06 Achiam et al. (2023), Qwen2.5 series Yang et al. (2024a), Internlm2.5 series Cai et al. (2024), Llama3.1 series Dubey et al. (2024), MiniCPM3-4B Hu et al. (2024), as well as 6 medical large language models, comprising the Apollo series Zheng et al. (2025); Wang et al. (2024c), and HuatuoGPT2 series Chen et al. (2024). We omit the "chat" and "instruct" in Table 2.
@@
Methods. We assess six prompt-based methodologies: (1) Standard Prompt, utilizing solely the original rules and patient notes; (2) Zero-Shot CoT Kojima et al. (2022), enhanced with the directive "Let’s think step by step"; (3) For CF-IR prompt, we leverage human-reviewd fact templates and logical rules to derive the answer; (4) Specifically, we examine the One-shot CoT Wei et al. (2022); (5) To explore the ability of LLM for rule representation enhancement, we introduce ACF-IR, an automated framework that enables LLMs to decompose rules and then conduct CF-IR; (6) Additonally, we set One-Shot CF-IR.
For each indicator, we selected an example outside of CMQCIC-Bench. The outputs for these examples were generated by GPT-4o in a zero-shot setting and were carefully reviewed and annotated by human evaluators.
@@ 5.2 Main Results
6 Empirical Analysis and Discussion
@@ 6.1 Error Analysis
Building on prior work Khandekar et al. (2025), we employ DeepSeek to classify error types by comparing LLM outputs with ground truth in CMQCIC-Bench, facilitating a granular error analysis across LLMs. Since incorrect clinical facts can propagate and affect inferential rule reasoning, we focus on identifying the earliest error type. A manual review of 200 randomly sampled DeepSeek-annotated errors confirmed an 87% accuracy, validating our approach for analyzing error types in all CF-IR responses. As shown in Table 4, providing demonstrations reduces Type A and B errors, highlighting the value of exemplars. While CF-IR does not mitigate clinical fact verification errors, it significantly improves reasoning accuracy due to its structured logical framework. Further details are available in Appendix E.1 (Tables 11 and 12).
@@ 6.3 Analysis on Clinical Fact Verification and Inferential Rule Abilities
While the CF-IR method enhances inference performance, we further investigate its two core capabilities: Clinical Fact Verification and Inferential Rule Reasoning. For Clinical Fact Verification, we define the input as <Patient Note, Templated Clinical Fact, Question> and the output as <Reasoning, Final Answer>, evaluated using Fact Faithfulness and Fact Correctness. Unlike step-wise reasoning, we test each fact independently to avoid contextual interference. For Inferential Rule Reasoning, providing verified facts as input to minimize errors, the input is <Verified Clinical Facts, Logical Rules>, and the output is <Explanation, Final Answer>, evaluated using labels like ’True/Yes’ or ’False/No’ for both natural (NL) and symbolic (SY) languages. All experiments are conducted in a zero-shot setting. Additional results are available in Table 13 in Appendix E.2.
As shown in Figures 5, we find that: (1) Both Qwen and Apollo exhibit performance-scale correlations across capabilities. (2) Fact verification performance significantly declines, consistent with Table 5. (3) For inferential reasoning, Qwen performs comparably in natural and symbolic settings, while Apollo shows stronger natural language robustness. (4) With correct facts, Qwen surpasses previous best results (standard, CoT, CF-IR) in zero-shot settings, whereas Apollo underperforms, likely due to Qwen’s extensive logical reasoning training. See Appendix E.2 for additional results.
@@ 6.4 Benefit of fine-tuning
We fine-tuned the Qwen2.5-3B-Instruct model using LoRA Hu et al. (2021) for 3 epochs on the CMQCIC-Bench, with evaluation on CMQCIC-Private. The data format follows ACF-IR: Input: <Instruction, Patient Note, Rule>; Output: <Knowledge, Templated Clinical Facts, Logical Rules, Clinical Fact Verification, Inferential Rule Reasoning, Final Answer>. As depicted in Figure 6, the fine-tuned 3B model achieves comparable or superior performance to the 7B model in real-world scenarios. In zero-shot settings, it demonstrates significant improvements across all methods, with gains of 6.7, 8.3, 14.3, and 13.7, confirming the feasibility of distillation rule enhancement for smaller models. More details are provided in Appendix B.
@@ 7.1 Rule-based LLM reasoning
While reasoning demonstrated a fundamental capability of LLM on applications Li et al. (2024), there are many research such as CoT Wei et al. (2022), CoT-sc Wang et al. (2023b), ToT Yao et al. (2024), etc. However, there are more attention on rule-enhanced methods Yang et al. (2023b); Sun et al. (2023); Mu et al. (2023). Reasoning based on facts and deriving answers from logical rules is referred to as inferential rule following ability Sun et al. (2024). Leveraging such ability that integrating explicit rules with LLMs has gained significant attention. For instance, Servantez et al. (2024b) utilized the IRAC framework to tackle legal tasks with LLMs, emphasizing the application of legal rules. Additionally, Wang et al. (2024b) proposed a neurosymbolic framework for multi-step rule application. Despite the current limitations of LLMs in rule-based reasoning Yang et al. (2024c), our work demonstrates that such rule-based reasoning outperforms CoT reasoning in the MQCIC task.
@@ 8 Conclusion
Acknowledgments
This paper was supported by the National Natural Science Foundation of China (No. 62306112), Shanghai Sailing Program (No. 23YF1409400), and Shanghai Pilot Program for Basic Research (No. 22TQ1400100-20).
Appendix A Models
@@
Appendix B Training Details
Appendix C Discussion on Test-Time Scaling
While test-time scaling like (CoT-SC, ToT or O1, R1) has attracted significant research attention, practical deployment in clinical settings requires careful consideration of GPU resource constraints, as hospitals typically have limited computational capacity. To address this, we quantitatively compare the inference efficiency of our method against CoT in Table 6, with measurements reported in GPU-hours.
@@ Appendix C Discussion on Test-Time Scaling
| Models | CoT | CF-IR |
| DeepSeek-V3 | 3h40min | 3h29min |
| DeepSeek-R1 | 10h10min | 9h30min |
Comparing inference times between chat and reasoning models in Table 7, we observe that while the reasoning model demonstrates better scalability, the performance gains remain marginal in Table 8. In the zero-shot setting, R1 did not demonstrate significant improvement. In contrast, our CF-IR framework outperformed the long cot method.
| Standard | CoT | CF-IR | |
| Qwen2.5-7B | 66.49 | 82.80 | 82.92 |
| Qwen2.5-14B | 78.79 | 86.49 | 87.21 |
| Qwen2.5-32B | 75.54 | 82.03 | 86.11 |
| Qwen2.5-72B | 87.77 | 87.51 | 92.73 |
| llama3.1-70B | 82.54 | 85.85 | 85.47 |
| GPT-4o | 77.45 | 88.91 | 91.84 |
| DeepSeek-V3 | 82.67 | 86.83 | 91.84 |
| DeepSeek-R1 | 81.01 | 82.99 | 92.73 |
@@
Appendix D Prompt
Here are the zero-shot prompt templates for the data construction, rule representation enhancement, clinical fact-based inferential rule reasoning method, and other prompt-based methods we used in this paper.
For each indicator rule question, we provide a list of facts that should be extracted. The prompt for clinical fact extraction is shown in Figure 7. The main fields are in the Table 9 and Table 10.
Prompt for Rule Enhancement. As mentioned in Section 4, we leverage LLMs to transform rules through three steps: Knowledge Enhancement, Rule Decomposition, and Clinical Fact Templazation. As shown in Figure 8, 9 and 10, we leverage GPT-4o first to generate the relative knowledge, logical rules, and templated clinical facts separately. Prompt for Different Methods. The detail prompt of different method as shown in Figure 11, Figure 12 and Figure 13.
Appendix E Additional Analysis Details
E.1 Error Analysis Details
As shown in Table 11 and Table 12 we observed: (1) Although the one-shot method reduces errors across both types in both approaches, the CoT method still results in more Type B errors, which may be due to the differing reasoning paths in the examples. (2) The CF-IR method effectively reduces Type B errors, but when it comes to Type A errors, the issue seems to be more related to the model’s intrinsic capabilities, which our method has not been able to enhance or activate effectively.
E.2 Clinical Fact Verification and Inferential Rule